# High-order persistence of resonant caustics in perturbed circular billiards

### v1

@@ 1. Introduction
A caustic is a curve such that any billiard trajectory, once tangent to the curve, stays tangent after every reflection. The robustness of a convex caustic is closely related to the arithmetic properties of its rotation number $\rho\in(0,1)$, a number that measures the number of turns around the caustic per impact. Tangent lines to the caustic can be counterclockwise or clockwise oriented. We fix the counterclockwise orientation, so $\rho\in(0,1/2]$. Lazutkin [33] showed that for any smooth strictly convex domain there is a positive measure Cantor set $\mathcal{R}\subset(0,1/2)$ of Diophantine rotation numbers that accumulates to 00 such that there is a caustic for any rotation number $\rho\in\mathcal{R}$. These caustics persist under smooth deformations of the domain [40].
@@
Fourier coefficients of $2\pi$-periodic functions are denoted with a hat: $a(t)=\sum_{l\in\mathbb{Z}}\hat{a}_{l}{\rm e}^{\mskip 2.0mu{\rm i}\mskip 1.0mult}$. Given a $2\pi$-periodic smooth function $a(t)$ and a subset $R\subset\mathbb{Z}$, let $\mu_{R}\{a(t)\}=\sum_{l\in R}\hat{a}_{l}{\rm e}^{\mskip 2.0mu{\rm i}\mskip 1.0mult}$ be the projection of $a(t)$ onto its $R$-harmonics. We only consider the cases $R=q\mathbb{Z}$ or $R=q\mathbb{Z}^{\star}$ with =$\mathbb{Z}^{\star}=\mathbb{Z}\setminus\{0\}$.
@@
[…]
@@
The $\mathop{\rm O}\nolimits(\epsilon)$-persistence result in Theorem 1 is just a reformulation of the main theorem in [41]. Condition $\mu_{q\mathbb{Z}^{\star}}\{h_{m}+\zeta_{m}\}=0$ is equivalent $\mu_{q\mathbb{Z}}\{h^{\prime}_{m}+\zeta^{\prime}_{m}\}=0$. In particular, condition $\mu_{q\mathbb{Z}^{\star}}\{h_{2}+\theta_{1}^{2}/2\}$ is equivalent to $\mu_{q\mathbb{Z}}\{h^{\prime}_{2}+\theta_{1}\theta^{\prime}_{1}\}=0$.
@@
$\chi=\chi(\Gamma_{\epsilon},q)=\begin{cases}1+2\left\lceil(q-n)/2n\right\rceil,&\text{for anti-centrally symmetric $\Gamma_{\epsilon}$ and odd $q$,}\\ 2\lceil q/2n\rceil,&\mbox{for anti-centrally symmetric $\Gamma_{\epsilon}$ andeven $q$},\\ \lceil 2q/n\rceil,&\mbox{for centrally symmetric $\Gamma_{\epsilon}$ and odd $q$},\\ \lceil q/n\rceil,&\mbox{otherwise.}\end{cases}$
@@
The map $q\mapsto\chi(\Gamma_{\epsilon},q)$ is unbounded for any polynomial deformation $\Gamma_{\epsilon}$ of degree $\leq n$, since $\chi(\Gamma_{\epsilon},q)\asymp 2q/n$ as odd $q\to+\infty$ for centrally symmetric deformations, and $\chi(\Gamma_{\epsilon},q)\asymp q/n$ as $q\to+\infty$ otherwise. The experiments described in [37, Numerical Result 5], which cover degrees $3\leq n\leq 8$ and periods $3\leq q\leq 100$, suggest that none of the $p/q$-resonant caustics $\mathop{\rm O}\nolimits(\epsilon^{\chi})$-persists under monomial deformations (4) with $P(x,y;\epsilon)=1-\epsilon y^{n}$ and $n\geq 3$. Its proof requires to check that $\mu_{q\mathbb{Z}^{\ast}}\{h_{\chi}+\zeta_{\chi}\}\neq 0$, which is a challenge. If it were true, that monomial deformations would break all resonant caustics in such a way that there would be breakups of any order, because the map $q\mapsto\chi(\Gamma_{\epsilon},q)\in\mathbb{N}$ is exhaustive.
@@ 2. Existence of smooth convex resonant caustics
[…]
@@
where $\psi_{j}=\sigma\{\varphi_{j}\}/2$ and $\theta_{j}=\delta\{\varphi_{j}\}/2$. Periodicities $\varphi_{j+q}=\varphi_{j}+2\pi p$, $\psi_{j+q}=\psi_{j}+2\pi p$ and $\theta_{j+q}=\theta_{j}$ imply that $\mu\{S(\varphi_{j_{0}+j},\varphi_{j_{0}+j+1})\}=2\mu\{h(\psi_{j_{0}+j})\sin\theta_{j_{0}+j}\}=A\{\varphi_{j}\}/q$ for all $j_{0}\in\mathbb{Z}$. Critical points of the action, which we call $p/q$-periodic configurations, can be lifted to full $p/q$-periodic orbits of the billiard map $f$ by taking $\lambda_{j}=-\partial_{1}S(\varphi_{j},\varphi_{j+1})=\partial_{2}S(\varphi_{j-1},\varphi_{j})$. Thus, any $p/q$-periodic configuration defines a $p/q$-periodic billiard trajectory inside $\Gamma$ with side angles $\varphi_{j}$, normal angles $\psi_{j}=\sigma\{\varphi_{j}\}/2$ and incidence-reflection angles $\theta_{j}=\delta\{\varphi_{j}\}/2$. We characterize such configurations in the next proposition, where we also recall a formula for the length of their corresponding periodic billiard trajectories.
@@ A $p/q$-periodic sequence $\{\varphi_{j}\}$ is a critical point of the action if and only if
We look for necessary conditions for the existence of convex resonant caustics inside $\Gamma$. Tangent lines to such a caustic can be counterclockwise or clockwise oriented. We fix the counterclockwise orientation, so we assume that $p/q<1/2$ from now on. Any $p/q$-resonant convex caustic gives rise to a 1-parameter family of $p/q$-periodic orbits of the billiard map. We want to parametrize this family using a dynamical parameter $t\in\mathbb{R}$ in which the billiard map acts as the constant shift $t\mapsto t+\omega$ with angular frequency $\omega=2\pi p/q$. The dynamical parameter is not unique. If $a(t)$ is any smooth $\omega$-periodic function such that $01+a^{\prime}(t)>$ 0, then $s=t+a(t)$ is another dynamical parameter.
@@
    where $\tilde{f}:\tilde{\mathbb{A}}_{\Gamma}\to\tilde{\mathbb{A}}_{\Gamma}$, $\tilde{g}:\mathbb{R}\to\mathbb{R}$ and $\tilde{r}:\mathbb{R}\to\mathbb{R}$ are the corresponding lifts. Identity $\tilde{r}^{q}(\varphi)=\varphi+q\omega$ is the key. It follows from the definition of $p/q$-resonant caustic. It implies that $\varphi\mapsto\tilde{r}(\varphi)$ becomes the constant shift $t\mapsto t+\omega$ in the smooth parameter
@@ 3. Persistence of smooth convex resonant caustics
$\partial_{2}S_{\epsilon}(\varphi_{\epsilon}(t-\omega),\varphi_{\epsilon}(t))+\partial_{1}S_{\epsilon}(\varphi_{\epsilon}(t),\varphi_{\epsilon}(t+\omega))=\mathop{\rm O}\nolimits(\epsilon^{m+1})\quad\mbox{ as }\epsilon\to 0,$
which stands out its Lagrangian nature. It was inspired by the proofs of the existence of (resonant and nonresonant) rotational invariant curves contained in [34, 31]. Condition (A)m follows the variational approach in [17].
Remark 4.
  1. (G)m
It means that $f^{q}_{\epsilon}$ projects $G^{\bullet}_{\epsilon}=\mathop{\rm graph}\nolimits(g^{\bullet}_{\epsilon})$ onto $G^{\star}_{\epsilon}=\mathop{\rm graph}\nolimits(g^{\star}_{\epsilon})$ in the vertical direction. None of these two graphs have to coincide with $G_{\epsilon}=c_{\epsilon}(\mathbb{T})$, but all of them are $\mathop{\rm O}\nolimits(\epsilon^{m+1})$-close.
@@ 4. High-order persistence in deformed circles
Besides, $\tilde{\mathcal{Q}}_{1}=0$, ~$\tilde{\mathcal{R}}_{1}=ch_{1}$, $\tilde{\mathcal{Q}}_{2}=sh^{\prime\prime}_{1}\psi_{1}+ch^{\prime}_{1}\theta_{1}$ and ~$\tilde{\mathcal{R}}_{2}=ch_{2}+ch^{\prime}_{1}\psi_{1}-sh_{1}\theta_{1}-c\theta^{2}_{1}/2$.
The proof of Lemma 7 is postponed to Appendix B. The first coefficients are obtained from the explicit recurrences for $\tilde{\mathcal{Q}}_{k}$ and ~$\tilde{\mathcal{R}}_{k}$ given in Lemma 15.
@@
Assume that the $p/q$-resonant caustic $\mathop{\rm O}\nolimits(\epsilon^{k-1})$-persists for some order $k\in\mathbb{N}$.
@@
[…]
@@
[…]
@@ 5. Polynomial deformations of the unit circle
    $\Gamma_{\epsilon}$ is anti-centrally symmetric, $q\neq m\mod 2$, $m\geq 2$ and $q>n(m-1)$.
@@
  • If $\Gamma_{\epsilon}$ is anti-centrally symmetric, then $nm<q$ or $n(m-1)<q$ with $m\neq q\mod 2$ and $m\geq 2$;
@@ Acknowledgments
C. E. K. gratefully acknowledges support from the European Research Council (ERC) through the Advanced Grant “SPERIG” (#885 707) . R. R.-R. was supported in part by the grant PID-2021-122954NB-100 which was funded by MCIN/AEI/10.13039/501100011033 and “ERDF: A way of making Europe”. R. R.-R. thanks to Pau Martín and Tere Seara for many long and stimulating conversations on related problems.
@@ Appendix A Shift, sum, difference and average operators
has a smooth $2\pi$-periodic solution $a(t)$ if and only if $\mu\{b(t)\}=0$, in which case the solution is unique under the additional condition $\mu\{a(t)\}=0$. Analogously, equation
@@
has a smooth $2\pi$-periodic solution $\tilde{a}(t)$ if and only if $\mu\{b(t)\}\equiv\omega$, in which case the solution is unique under the additional condition $\mu\{\tilde{a}(t)\}=0$.
@@
[…]
$\hat{a}_{l}=\hat{b}_{l}/({\rm e}^{\mskip 2.0mu{\rm i}\mskip 1.0mul\omega}-1),\qquad\forall l\not\in q\mathbb{Z},$
[…]

### v2

@@ 1. Introduction
A caustic is a curve such that any billiard trajectory, once tangent to the curve, stays tangent after every reflection. The robustness of a convex caustic is closely related to the arithmetic properties of its rotation number $\rho\in(0,1)$, a number that measures the number of turns around the caustic per impact. Tangent lines to the caustic can be counterclockwise or clockwise oriented. We fix the counterclockwise orientation, so $\rho\in(0,1/2]$. Lazutkin [33] showed that for any smooth strictly convex domain there is a positive measure Cantor set $\mathcal{R}\subset(0,1/2)$ of Diophantine rotation numbers that accumulates to 0 such that there is a caustic for any rotation number $\rho\in\mathcal{R}$. These caustics persist under smooth deformations of the domain [40].
@@
Fourier coefficients of $2\pi$-periodic functions are denoted with a hat: $a(t)=\sum_{l\in\mathbb{Z}}\hat{a}_{l}{\rm e}^{\mskip 2.0mu{\rm i}\mskip 1.0mult}$. Given a $2\pi$-periodic smooth function $a(t)$ and a subset $R\subset\mathbb{Z}$, let $\mu_{R}\{a(t)\}=\sum_{l\in R}\hat{a}_{l}{\rm e}^{\mskip 2.0mu{\rm i}\mskip 1.0mult}$ be the projection of $a(t)$ onto its $R$-harmonics. We only consider the cases $R=q\mathbb{Z}$ or $R=q\mathbb{Z}^{\star}$ with $\mathbb{Z}^{\star}=\mathbb{Z}\setminus\{0\}$.
@@
    $\theta_{1}(t)=\sum_{l\not\in q\mathbb{Z}\cup\{-1,1\}}\nu_{l}(\rho)\hat{h}_{1,l}\ {\rm e}^{\mskip 2.0mu{\rm i}\mskip 1.0mult}\quad\mbox{ if }\quad h_{1}(\psi)=\sum_{l\in\mathbb{Z}}\hat{h}_{1,l}\ {\rm e}^{\mskip 2.0mu{\rm i}\mskip 1.0mul\psi}.$
@@
The $\mathop{\rm O}\nolimits(\epsilon)$-persistence result in Theorem 1 is just a reformulation of the main theorem in [41]. Condition $\mu_{q\mathbb{Z}^{\star}}\{h_{m}+\zeta_{m}\}=0$ is equivalent to $\mu_{q\mathbb{Z}}\{h^{\prime}_{m}+\zeta^{\prime}_{m}\}=0$. In particular, condition $\mu_{q\mathbb{Z}^{\star}}\{h_{2}+\theta_{1}^{2}/2\}$ is equivalent to $\mu_{q\mathbb{Z}}\{h^{\prime}_{2}+\theta_{1}\theta^{\prime}_{1}\}=0$.
@@
[…]
@@
The map $q\mapsto\chi(\Gamma_{\epsilon},q)$ is unbounded for any polynomial deformation $\Gamma_{\epsilon}$ of degree $\leq n$, since $\chi(\Gamma_{\epsilon},q)\asymp 2q/n$ as odd $q\to+\infty$ for centrally symmetric deformations, and $\chi(\Gamma_{\epsilon},q)\asymp q/n$ as $q\to+\infty$ otherwise. The experiments described in [37, Numerical Result 5], which cover degrees $3\leq n\leq 8$ and periods $3\leq q\leq 100$, suggest that none of the $p/q$-resonant caustics $\mathop{\rm O}\nolimits(\epsilon^{\chi})$-persists under monomial deformations (4) with $P(x,y;\epsilon)=1-\epsilon y^{n}$ and $n\geq 3$. Its proof requires to check that $\mu_{q\mathbb{Z}^{\ast}}\{h_{\chi}+\zeta_{\chi}\}\neq 0$, which is a challenge. If $\Gamma_{\epsilon}$ is a polynomial deformation of the unit circle of degree $\leq n$ such that $\mu_{q\mathbb{Z}^{\ast}}\{h_{\chi}+\zeta_{\chi}\}\neq 0$ for all $p/q\in(0,1/2)$, then $\Gamma_{\epsilon}$ would break all resonant caustics in such a way that there would be breakups of any order, because the map $q\mapsto\chi(\Gamma_{\epsilon},q)\in\mathbb{N}$ is exhaustive.
@@ 2. Existence of smooth convex resonant caustics
$z:\mathbb{T}\to\Gamma\subset\mathbb{R}^{2}\simeq\mathbb{C},\qquad z(\varphi)=\big{(}x(\varphi),y(\varphi)\big{)}=h(\varphi){\rm e}^{\mskip 2.0mu{\rm i}\mskip 1.0mu\varphi}+h^{\prime}(\varphi){\rm e}^{\mskip 2.0mu{\rm i}\mskip 1.0mu(\varphi+\pi/2)},$
@@
where $\psi_{j}=\sigma\{\varphi_{j}\}/2$ and $\theta_{j}=\delta\{\varphi_{j}\}/2$. Periodicities $\varphi_{j+q}=\varphi_{j}+2\pi p$, $\psi_{j+q}=\psi_{j}+2\pi p$ and $\theta_{j+q}=\theta_{j}$ imply that $\mu\{S(\varphi_{j_{0}+j},\varphi_{j_{0}+j+1})\}=2\mu\{h(\psi_{j_{0}+j})\sin\theta_{j_{0}+j}\}=A\{\varphi_{j}\}/q$ for all $j_{0}\in\mathbb{Z}$. Critical points of the action, which we call $p/q$-periodic configurations, can be lifted to full $p/q$-periodic orbits of the billiard map $f$ by taking $\lambda_{j}=-\partial_{1}S(\varphi_{j},\varphi_{j+1})=\partial_{2}S(\varphi_{j-1},\varphi_{j})$. Thus, any $p/q$-periodic configuration defines a $p/q$-periodic billiard trajectory inside $\Gamma$ with side angles $\varphi_{j}$, normal angles $\psi_{j}=\sigma\{\varphi_{j}\}/2$ and incidence-reflection angles $\theta_{j}=\delta\{\varphi_{j}\}/2$.
We characterize such configurations in the next proposition, where we also recall a formula for the length of their corresponding periodic billiard trajectories.
@@ A $p/q$-periodic sequence $\{\varphi_{j}\}$ is a critical point of the action if and only if
We look for necessary conditions for the existence of convex resonant caustics inside $\Gamma$. Tangent lines to such a caustic can be counterclockwise or clockwise oriented. We fix the counterclockwise orientation, so we assume that $p/q<1/2$ from now on. Any $p/q$-resonant convex caustic gives rise to a 1-parameter family of $p/q$-periodic orbits of the billiard map. We want to parametrize this family using a dynamical parameter $t\in\mathbb{R}$ in which the billiard map acts as the constant shift $t\mapsto t+\omega$ with angular frequency $\omega=2\pi p/q$. The dynamical parameter is not unique. If $a(t)$ is any smooth $\omega$-periodic function such that $1+a^{\prime}(t)>0$, then $s=t+a(t)$ is another dynamical parameter.
@@
    where $\tilde{f}:\tilde{\mathbb{A}}_{\Gamma}\to\tilde{\mathbb{A}}_{\Gamma}$, $\tilde{g}$$\mathbb{R}\to\mathbb{R}$ and $\tilde{r}$$\mathbb{R}\to\mathbb{R}$ are the corresponding lifts. Identity $\tilde{r}^{q}(\varphi)=\varphi+q\omega$ is the key. It follows from the definition of $p/q$-resonant caustic. It implies that $\varphi\mapsto\tilde{r}(\varphi)$ becomes the constant shift $t\mapsto t+\omega$ in the smooth parameter
@@ 3. Persistence of smooth convex resonant caustics
$\partial_{2}S_{\epsilon}(\varphi_{\epsilon}(t-\omega),\varphi_{\epsilon}(t))+\partial_{1}S_{\epsilon}(\varphi_{\epsilon}(t),\varphi_{\epsilon}(t+\omega))=\mathop{\rm O}\nolimits(\epsilon^{m+1})\quad\mbox{ as }\epsilon\to 0.$
Thus, Condition (F)m is related to the Lagrangian formulation as a second-order difference equation of the invariance condition for (nonresonant and resonant) rotational curves of exact twist maps. It was inspired by ideas contained in [34, 31]. Compare with [34, Eq. (3)] in the nonresonant setting and with [31, Eq. (1)] in the resonant setting. Finally, Condition (A)m follows the variational approach in [17].
@@
Remark 5.
  1. (G)m
It means that $f^{q}_{\epsilon}$ projects $G^{\bullet}_{\epsilon}=\mathop{\rm graph}\nolimits(g^{\bullet}_{\epsilon})$ onto $G^{\star}_{\epsilon}=\mathop{\rm graph}\nolimits(g^{\star}_{\epsilon})$ in the vertical direction. None of these two graphs have to coincide with $G_{\epsilon}=c_{\epsilon}(\mathbb{T})$, but all of them are $\mathop{\rm O}\nolimits(\epsilon^{m+1})$-close. Conditions (F)m or (P)m imply Condition (G)m. The proof of this implication is based on two ideas. First, we check that the restriction of the power map $f^{q}_{\epsilon}$ to the graph $G_{\epsilon}$ is $\mathop{\rm O}\nolimits(\epsilon^{m+1})$-close to the identity, although the error in the horizontal direction may be nonzero. Next, we use the twist condition to cancel the horizontal error, so we get the graphs $G^{\bullet}_{\epsilon},G^{\star}_{\epsilon}=G_{\epsilon}+\mathop{\rm O}\nolimits(\epsilon^{m+1})$. We omit the technical details since we do not need this result. On the contrary, we do not know whether Condition (G)m implies Conditions (F)m and (P)m, because it is not clear how to extract information about a single iterate of the perturbed map from information about the power map $f^{q}_{\epsilon}$.
@@ 4. High-order persistence in deformed circles
Besides, $\tilde{\mathcal{Q}}_{1}=0$, $\tilde{\mathcal{R}}_{1}=ch_{1}$, $\tilde{\mathcal{Q}}_{2}=sh^{\prime\prime}_{1}\psi_{1}+ch^{\prime}_{1}\theta_{1}$ and $\tilde{\mathcal{R}}_{2}=ch_{2}+ch^{\prime}_{1}\psi_{1}-sh_{1}\theta_{1}-c\theta^{2}_{1}/2$.
The proof of Lemma 7 is postponed to Appendix B. The first coefficients are obtained from the explicit recurrences for $\tilde{\mathcal{Q}}_{k}$ and $\tilde{\mathcal{R}}_{k}$ given in Lemma 15.
@@
Let $k\in\mathbb{N}$ be a fixed order. If the $p/q$-resonant caustic $\mathop{\rm O}\nolimits(\epsilon^{k-1})$-persists, then the following properties hold.
@@
We denote the Fourier coefficients of smooth $2\pi$-periodic functions with a hat, so we write $h_{1}(t)=\sum_{l\in\mathbb{Z}}\hat{h}_{1,l}{\rm e}^{\mskip 2.0mu{\rm i}\mskip 1.0mult}$ and $\theta_{1}(t)=\sum_{l\in\mathbb{Z}}\hat{\theta}_{1,l}{\rm e}^{\mskip 2.0mu{\rm i}\mskip 1.0mult}$.
@@
    $se^{-}_{l}\hat{\theta}_{1,l}=\big{(}ce^{-}_{l}-\mskip 2.0mu{\rm i}\mskip 1.0mulse^{+}_{l}\big{)}\hat{h}_{1,l},\qquad\forall l\not\in q\mathbb{Z}.$
@@ 5. Polynomial deformations of the unit circle
    $\Gamma_{\epsilon}$ is anti-centrally symmetric, $q\neq m\mod 2$, $m\geq 2$ and q>n(m1)q>n(m-1).
@@
  • If $\Gamma_{\epsilon}$ is anti-centrally symmetric, then $nm<q$ or n(m1)<qn(m-1)<q with $m\neq q\mod 2$ and $m\geq 2$;
@@ Acknowledgments
C. E. K. gratefully acknowledges support from the European Research Council (ERC) through the Advanced Grant “SPERIG” (#885 707) . R. R.-R. was supported in part by the grant PID-2021-122954NB-100 which was funded by MCIN/AEI/10.13039/501100011033 and “ERDF: A way of making Europe”. R. R.-R. thanks to Pau Martín and Tere Seara for many long and stimulating conversations on related problems. We are also indebted to the referee for suggestions that helped us to improve the exposition.
@@ Appendix A Shift, sum, difference and average operators
has a smooth $2\pi$-periodic solution $a(t)$ if and only if $\mu\{b(t)\}=0$, in which case the solution is unique once we fix its smooth $2\pi/q$-periodic average $\mu\{a(t)\}$. Analogously, equation
@@
has a smooth $2\pi$-periodic solution $\tilde{a}(t)$ if and only if $\mu\{b(t)\}\equiv\omega$, in which case the solution is unique once we fix its smooth $2\pi/q$-periodic average $\mu\{\tilde{a}(t)\}$.
@@
Conversely, if $\mu\{b(t)\}=0$, then $b(t)=\sum_{l\not\in q\mathbb{Z}}\hat{b}_{l}{\rm e}^{\mskip 2.0mu{\rm i}\mskip 1.0mult}$ so that
$\hat{a}_{l}=\hat{b}_{l}/({\rm e}^{\mskip 2.0mu{\rm i}\mskip 1.0mul\omega}-1),\qquad\forall l\not\in q\mathbb{Z}.$
defines the Fourier coefficients of the unique smooth $2\pi$-periodic solution $a(t)=\sum_{l\not\in q\mathbb{Z}}\hat{a}_{l}{\rm e}^{\mskip 2.0mu{\rm i}\mskip 1.0mult}$ such that $\mu\{a(t)\}=0$. Smoothness follows from the bounds $|\hat{a}|_{l}\leq\nu|\hat{b}_{l}|$ for all $l\not\in q\mathbb{Z}$, where $\nu=1/|{\rm e}^{\mskip 2.0mu{\rm i}\mskip 1.0mu2\pi/q}-1|>0$. Since $\delta\circ\mu=0$, we deduce that the average $\mu\{a(t)\}$ does not need to be zero, but can be determined freely.